OpenStudy (anonymous):

find the derivative: 1/ sqrt(x)

OpenStudy (anonymous):

\[1/\sqrt{x}\]

ganeshie8 (ganeshie8):

is it find, or deriving using first principle ? finding is easy \(\frac{1}{\sqrt{x}}\) = \(\frac{1}{x^\frac{1}{2}}\) = \(x^\frac{-1}{2}\)

OpenStudy (anonymous):

no derive, im trying to use quotent rule

ganeshie8 (ganeshie8):

ohhk...

ganeshie8 (ganeshie8):

can become tricky.... il give a try. others pls feel free to reply :D

OpenStudy (anonymous):

ive gotten \[-0.5\sqrt{x} - \sqrt{x} / x\]

OpenStudy (anonymous):

im not really shore how to simplify that

hartnn (hartnn):

 i have written root x as x ^ 1/2